OpenStudy (anonymous):

OK GUYS new question :) (easier, i think :P ) f(x) = 2x + 1 f(g(x)) = 1/(2x+1) What is g(x) ??

OpenStudy (amistre64):

how do we turn f(x) into f(g(x)) by altering the x part?

OpenStudy (amistre64):

it might be easier to say f(a) instead of f(x) limits the amount of x on x confusion to me

OpenStudy (anonymous):

but the point of using x, is that in the equation the variables are called x :P

OpenStudy (amistre64):

i know, but to separate the x in f and the x in g is useful .... say f(a) = 2a+1; then g(x) = a

OpenStudy (amistre64):

2a+1 = 1/(2x+1) solve for a, since a = g(x)

OpenStudy (amistre64):

or ... to keep your preciuous xs :) 2g(x) + 1 = 1/(2x+1) and solve for g(x)

OpenStudy (amistre64):

subtract 1 and divide off the 2

OpenStudy (anonymous):

HAWW I DIDN'T THINK OF THAT WAIT.

OpenStudy (anonymous):

so i got (2x + 1)/-x now what :P

OpenStudy (amistre64):

2g(x) + 1 = 1/(2x+1) 2g(x) = 1/(2x+1) - 1 g(x) = 1/2(2x+1) - 1/2 g(x) = 1/2(2x+1) - 1/2 is what i get

OpenStudy (anonymous):